Write the slope-intercept equation for the line that passes through (-10,5) and has slope 9/10.
Thank you in advance. I am just learning this with your help.

Answer by stanbon(75887)

Write the slope-intercept equation for the line that passes through (-10,5) and has slope 9/10.
----
Slope-intercept form is y = mx + b
where m is the slope.
---
Substitute for x,y, and m and solve for "b":
----
5 = (9/10)(-10)+b
---
5 = -9 + b
---
b = 14
=============
Equation
y = (9/10)x + 14
========================
Cheers,
Stan H.